# The deep multi-FBSDE method: a robust deep learning method for coupled FBSDEs

### v1

@@ Abstract
We introduce the deep multi-FBSDE method for robust approximation of coupled forward–backward stochastic differential equations (FBSDEs), focusing on cases where the deep BSDE method of Han, Jentzen, and E (2018) fails to converge. To overcome the convergence issues, we consider a family of FBSDEs that are equivalent to the original problem in the sense that they satisfy the same associated partial differential equation (PDE). Our algorithm proceeds in two phases: first, we approximate the initial condition for the FBSDE family, and second, we approximate the original FBSDE using the initial condition approximated in the first phase. Numerical experiments show that our method converges even when the standard deep BSDE method does not.
@@ 1 Introduction
  • A general purpose training strategy applicable general FBSDE systems, without restrictive assumptions such as the existence of an equivalent stochastic control problem, as in [1].
@@
The rest of this paper is organized as follows: Section 2 presents the FBSDE problem and its solution by the non-linear Feynann-Kac formula through the PDE. Section 3 introduces our proposed robust deep learning method, detailing its formulation and implementation. Numerical results demonstrating its advantages are presented in Section 4, and we end in Section 5 with conclusions and future directions.
@@ 2 Forward backward stochastic differential equations
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}b(s,X_{s},Y_{s},Z_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s})\text{d}W_{s},\\ Y_{t}=g(X_{T})+\int_{t}^{T}f(s,X_{s},Y_{s},Z_{s})\text{d}s-\int_{t}^{T}Z_{s}\sigma(s,X_{s})\text{d}W_{s},\quad t\in[0,T].\end{dcases}$ (1)
@@
$\begin{dcases}\frac{\partial v}{\partial t}(t,x)+\mathcal{L}v(t,x)+f\big{(}t,x,v(t,x),\mathrm{D}_{x}v(t,x)\sigma(t,x)\big{)}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(t,x)=g(x),&(t,x)\in\{T\}\times\mathbb{R}^{d}.\end{dcases}$ (2)
Here $\mathcal{L}$ is the infinitesimal generator, which when acting on a twice continuously differentiable function $\rho$ yields
@@
where $a=\sigma\sigma^{\top}$ and $\text{Tr}(\cdot)$ denotes the trace of a matrix. The connection between (1) and (2) is established through the nonlinear Feyman–Kac formula, see e.g., [32], which states that
@@ 3 The deep BSDE and deep multi-FBSDE methods
In this section, we first introduce the classical deep BSDE method and briefly discuss its convergence problems for coupled FBSDEs. Most importantly, we introduce our novel modification of the scheme. We use a linear quadratic Gaussian (LQG) problem to demonstrate both the convergence problems of the deep BSDE method and the convergence of our method. The details of the LQG problem can be found in Section 4.1.2 together complementary experimental results.
@@ 3.1 The deep BSDE method
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \mathbb{E}|Y_{T}^{y_{0},\zeta}-g(X_{T}^{y_{0},\zeta})|^{2},\quad\text{where}\\ X_{t}^{y_{0},\zeta}=x_{0}+\int_{0}^{t}b(s,X_{s}^{y_{0},\zeta},Y_{s}^{y_{0},\zeta},Z^{y_{0},\zeta}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta}=y_{0}-\int_{0}^{t}f(s,X_{s}^{y_{0},\zeta},Y_{s}^{y_{0},\zeta},Z^{y_{0},\zeta}_{s})\text{d}s+\int_{0}^{t}Z^{y_{0},\zeta}_{s}\sigma(s,X_{s}^{y_{0},\zeta}),\text{d}W_{s},\\ Z^{y_{0},\zeta}_{t}=\zeta(t,X_{t}^{y_{0},\zeta}),\quad t\in[0,T].\end{dcases}$ (3)
@@
$\begin{dcases}\underset{y_{0},\zeta_{0},\ldots,\zeta_{N-1}}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})\big{|}^{2},\quad\text{where}\\ X_{n+1}^{h,y_{0},\zeta}=X_{n}^{h,y_{0},\zeta}+b\big{(}t_{n},X_{n}^{h,y_{0},\zeta},Y_{n}^{h,y_{0},\zeta},Z_{n}^{h,y_{0},\zeta}\big{)}h+\sigma(s,X_{n}^{h,y_{0},\zeta})\Delta W_{n},\\ Y_{n+1}^{h,y_{0},\zeta}=Y_{n}^{h,y_{0},\zeta}-f\big{(}t_{n},X_{n}^{h,y_{0},\zeta},Y_{n}^{h,y_{0},\zeta},Z_{n}^{h,y_{0},\zeta}\big{)}h+Z_{n}^{h,y_{0},\zeta}\sigma\big{(}t_{n},X_{n}^{h,y_{0},\zeta}\big{)}\Delta W_{n},\\ Z_{n}^{h,y_{0},\zeta}=\zeta_{n}(X_{n}^{h,y_{0},\zeta}),\ Y_{0}^{h,y_{0},\zeta}=y_{0},\ X_{0}^{h,y_{0},\zeta}=x_{0},\quad n\in\{0,1,\ldots,N-1\}.\end{dcases}$ (4)
In the deep FBSDE method, we search for $y_{0}\in\mathbb{R}$, a simple trainable parameter, and for $n\in{0,\ldots,N-1}$, $\zeta_{n}$ in a parametric space defined by the architecture of a specified neural network.
@@
Under conditions that guarantee the same or similar a posteriori bound, such as those in [18], the function $\mathbb{E}|\hat{Y}_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})|^{2}$ is a highly suitable loss function and guarantees an approximation with an error bounded by the loss after training and the time step. When the assumptions are not met, the situation is less fortunate. Counterexamples of FBSDE from stochastic control are presented in [1] for which it is possible to make the right-hand side of (5) arbitrarily small, but with the left-hand side remaining large and discrete solutions converging to severely wrong processes. In this paper, we provide four more counterexamples. Thus, while the a posteriori bound in [18] is a powerful tool under its specific conditions, extrapolating it beyond these conditions is a shaky practice.
@@
[…]
This is the MSE with $y_{0}$ fixed and only $\zeta$ being optimized. If joint optimization of $y_{0}$ and $\zeta$, as in the deep BSDE method, is supposed to yield a good approximation of the FBSDE, then by necessity the minimum of $y_{0}\mapsto\mathrm{MSE}(y_{0})$ must be close to the true $Y_{0}$ of the FBSDE. For an LQG problem, detailed in Section 4.1.2 below, Figure 1 shows that the optimization landscape $y_{0}\mapsto\mathrm{MSE}(y_{0})$ instead has a minimum $y_{0}^{*}$ very far from the true $Y_{0}$. Figure 2 shows the exact $Y$ and the output of the deep FBSDE method, pathwise and in mean. We note that although the initial condition is significantly incorrect, the terminal condition is well approximated, i.e., $g(X_{N}^{h,y_{0},\zeta})-Y_{N}^{h,y_{0},\zeta}$ is small both on the pathwise level and in the mean-squared sense. It should be noted that the numerical algorithm has converged, as the loss function is no longer decreasing and $y_{0}$ remains constant (not shown here but Figure 6 qualitatively shows the same behavior for a different problem). Moreover, the algorithm is stable, with repeated runs yielding similar values for the loss function and $y_{0}$. The reader might wonder if a smaller time step $h$ in the approximation would improve approximation of $Y_{0}$. Figure 5 gives a negative answer to this valid thought by showing non-convergence (the case $K=1$ in that figure).
Figure 1: MSE plotted for different values of $y_{0}$ for the LQG-problem of Section 4.1.2.
@@
Figure 2: The almost analytical $Y$-process and the approxamtion obtained by using the deep FBSDE method. Left: One representative path. Right: A sample mean of size $2^{12}$.
@@ 3.2 The deep multi-FBSDE method
$\begin{dcases}X_{t}^{\psi}=x_{0}+\int_{0}^{t}\big{(}b(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})-\psi(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})\big{)}\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\psi})\text{d}W_{s},\\ Y_{t}^{\psi}=g(X_{T}^{\psi})+\int_{t}^{T}\big{(}f(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})-\big{\langle}\psi(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi}),Z_{s}^{\psi}\big{\rangle}\big{)}\text{d}s\\ \quad\quad\ -\int_{t}^{T}Z_{s}^{\psi}\sigma(s,X_{s}^{\psi})\text{d}W_{s},\quad t\in[0,T].\end{dcases}$ (7)
@@
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{T}^{y_{0},\zeta,\psi}-g(X_{T}^{y_{0},\zeta,\psi})\big{|}^{2},\quad\text{where}\\ X_{t}^{y_{0},\zeta,\psi}=x_{0}+\int_{0}^{t}b^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta,\psi}=y_{0}-\int_{0}^{t}f^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}Z^{y_{0},\zeta,\psi}_{s}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Z^{y_{0},\zeta,\psi}_{t}=\zeta(t,X_{t}^{y_{0},\zeta,\psi}),\quad t\in[0,T].\end{dcases}$ (8)
@@
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \sum_{k=1}^{K}\mathbb{E}\big{|}Y_{T}^{y_{0},\zeta,\psi_{k}}-g(X_{T}^{y_{0},\zeta,\psi_{k}})\big{|}^{2},\quad\text{where for }\psi\in\{\psi_{1},\dots,\psi_{K}\}\\ X_{t}^{y_{0},\zeta,\psi}=x_{0}+\int_{0}^{t}b^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta,\psi}=y_{0}-\int_{0}^{t}f^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}Z^{y_{0},\zeta,\psi}_{s}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Z^{y_{0},\zeta,\psi}_{t}=\zeta(t,X_{t}^{y_{0},\zeta,\psi}),\quad t\in[0,T].\end{dcases}$ (9)
@@
$\begin{dcases}\underset{y_{0},\zeta_{0},\ldots,\zeta_{N-1}}{\mathrm{minimize}}\ \sum_{k=1}^{K}\mathbb{E}\big{|}Y_{N}^{h,y_{0},\zeta,\psi_{k}}-g(X_{N}^{h,y_{0},\zeta,\psi_{k}})\big{|}^{2},\quad\text{where for }\psi\in\{\psi_{1},\dots,\psi_{K}\}\\ X_{n+1}^{h,y_{0},\zeta,\psi}=X_{n}^{h,y_{0},\zeta,\psi}+b^{\psi}\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi},Y_{n}^{h,y_{0},\zeta,\psi},Z_{n}^{h,y_{0},\zeta,\psi}\big{)}h+\sigma(s,X_{n}^{h,y_{0},\zeta,\psi})\Delta W_{n},\\ Y_{n+1}^{h,y_{0},\zeta,\psi}=Y_{n}^{h,y_{0},\zeta,\psi}-f^{\psi}\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi},Y_{n}^{h,y_{0},\zeta,\psi},Z_{n}^{h,y_{0},\zeta,\psi}\big{)}h+Z_{n}^{h,y_{0},\zeta,\psi}\sigma\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi}\big{)}\Delta W_{n},\\ Z_{n}^{h,y_{0},\zeta,\psi}=\zeta_{n}(X_{n}^{h,y_{0},\zeta,\psi}\big{)},\ X_{0}^{h,y_{0},\zeta,\psi}=x_{0},\ Y_{0}^{h,y_{0},\zeta,\psi}=y_{0},\\ \text{for }n\in\{0,1,\ldots,N-1\},\end{dcases}$ (10)
@@
$\begin{dcases}\underset{\zeta_{0},\ldots,\zeta_{N-1}}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{N}^{h,\zeta}-g(X_{N}^{h,\zeta})\big{|}^{2},\quad\text{where}\\ X_{n+1}^{h,\zeta}=X_{n}^{h,\zeta}+b\big{(}t_{n},X_{n}^{h},Y_{n}^{h,\zeta},Z_{n}^{h}\big{)}h+\sigma(s,X_{n}^{h,\zeta})\Delta W_{n},\\ Y_{n+1}^{h,\zeta}=Y_{n}^{h,\zeta}-f\big{(}t_{n},X_{n}^{h,\zeta},Y_{n}^{h,\zeta},Z_{n}^{h,\zeta}\big{)}h+Z_{n}^{h,\zeta}\sigma\big{(}t_{n},X_{n}^{h,\zeta}\big{)}\Delta W_{n},\\ Z_{n}^{h,\zeta}=\zeta_{n}\big{(}X_{n}^{h,\zeta}\big{)},\ Y_{0}^{h,\zeta}(m)=y_{0}^{*},\quad\text{for }n\in\{0,1,\ldots,N-1\},\end{dcases}$ (11)

### v2

@@ Abstract
We introduce the deep multi-FBSDE method for robust approximation of coupled forward–
backward stochastic differential equations (FBSDEs), focusing on cases where the deep BSDE method of Han, Jentzen, and E (2018) fails to converge. To overcome the convergence issues, we consider a family of FBSDEs that are equivalent to the original problem in the sense that they satisfy the same associated partial differential equation (PDE). Our algorithm proceeds in two phases: first, we approximate the initial condition for the FBSDE family, and second, we approximate the original FBSDE using the initial condition approximated in the first phase. Numerical experiments show that our method converges even when the standard deep BSDE method does not.
@@ 1 Introduction
  • A general purpose training strategy applicable to general FBSDE systems, without restrictive assumptions such as the existence of an equivalent stochastic control problem, as in [1].
@@
The rest of this paper is organized as follows: Section 2 presents the FBSDE problem and its solution by the non-linear Feynman–Kac formula through the PDE. Section 3 introduces our proposed robust deep learning method, detailing its formulation and implementation. Numerical results demonstrating its advantages are presented in Section 4, and we end in Section 5 with conclusions and future directions.
@@ 2 Forward backward stochastic differential equations
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}b(s,X_{s},Y_{s},Z_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s})\text{d}W_{s},\\ Y_{t}=g(X_{T})+\int_{t}^{T}f(s,X_{s},Y_{s},Z_{s})\text{d}s-\int_{t}^{T}\langle Z_{s},\sigma(s,X_{s})\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (1)
@@
$\begin{dcases}\frac{\partial v}{\partial t}(t,x)+\mathcal{L}v(t,x)+f\big{(}t,x,v(t,x),\mathrm{D}_{x}v(t,x)\sigma(t,x)\big{)}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(T,x)=g(x),&x\in\mathbb{R}^{d}.\end{dcases}$ (2)
Here $\mathcal{L}$ is the infinitesimal generator, acting on a twice continuously differentiable functions $\rho$ by:
@@
where $a=\sigma\sigma^{\top}$ and $\text{Tr}(\cdot)$ denotes the trace of a matrix. The connection between (1) and (2) is established through the nonlinear Feynman–Kac formula, see e.g., [32], which states that
@@ 3 The deep BSDE and deep multi-FBSDE methods
In this section, we first introduce the classical deep BSDE method and briefly discuss its convergence problems for coupled FBSDEs. Most importantly, we introduce our novel modification of the scheme. We use a Linear Quadratic (LQ) problem to demonstrate both the convergence problems of the deep BSDE method and the convergence of our method. The details of the LQ-problem can be found in Section 4.1.2 together complementary experimental results.
@@ 3.1 The deep BSDE method
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \mathbb{E}|Y_{T}^{y_{0},\zeta}-g(X_{T}^{y_{0},\zeta})|^{2},\quad\text{where}\\ X_{t}^{y_{0},\zeta}=x_{0}+\int_{0}^{t}b(s,X_{s}^{y_{0},\zeta},Y_{s}^{y_{0},\zeta},Z^{y_{0},\zeta}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta}=y_{0}-\int_{0}^{t}f(s,X_{s}^{y_{0},\zeta},Y_{s}^{y_{0},\zeta},Z^{y_{0},\zeta}_{s})\text{d}s+\int_{0}^{t}\langle Z^{y_{0},\zeta}_{s},\sigma(s,X_{s}^{y_{0},\zeta})\text{d}W_{s}\rangle,\\ Z^{y_{0},\zeta}_{t}=\zeta(t,X_{t}^{y_{0},\zeta}),\quad t\in[0,T].\end{dcases}$ (3)
@@
$\begin{dcases}\underset{y_{0},\zeta=(\zeta_{0},\ldots,\zeta_{N-1})}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})\big{|}^{2},\quad\text{where}\\ X_{n+1}^{h,y_{0},\zeta}=X_{n}^{h,y_{0},\zeta}+b\big{(}t_{n},X_{n}^{h,y_{0},\zeta},Y_{n}^{h,y_{0},\zeta},Z_{n}^{h,y_{0},\zeta}\big{)}h+\sigma(t_{n},X_{n}^{h,y_{0},\zeta})\Delta W_{n},\\ Y_{n+1}^{h,y_{0},\zeta}=Y_{n}^{h,y_{0},\zeta}-f\big{(}t_{n},X_{n}^{h,y_{0},\zeta},Y_{n}^{h,y_{0},\zeta},Z_{n}^{h,y_{0},\zeta}\big{)}h+\big{\langle}Z_{n}^{h,y_{0},\zeta},\sigma\big{(}t_{n},X_{n}^{h,y_{0},\zeta}\big{)}\Delta W_{n}\big{\rangle},\\ Z_{n}^{h,y_{0},\zeta}=\zeta_{n}(X_{n}^{h,y_{0},\zeta}),\ Y_{0}^{h,y_{0},\zeta}=y_{0},\ X_{0}^{h,y_{0},\zeta}=x_{0},\quad n\in\{0,1,\ldots,N-1\}.\end{dcases}$ (4)
In the deep BSDE method, we search for $y_{0}\in\mathbb{R}$, a simple trainable parameter, and for $\zeta=(\zeta_{n})_{n=0}^{N-1}$ in a parametric space defined by the architecture of a specified neural network.
@@
Under conditions that guarantee the same or similar a posteriori bound, such as those in [18], the function $\mathbb{E}|\hat{Y}_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})|^{2}$ is a highly suitable loss function and guarantees an approximation with an error bounded by the loss after training and the time step. When the assumptions are not met, the situation is less fortunate. Counterexamples of FBSDEs from stochastic control are presented in [1] for which it is possible to make the right-hand side of (5) arbitrarily small, but with the left-hand side remaining large and discrete solutions converging to severely wrong processes. In this paper, we provide four more counterexamples. Thus, while the a posteriori bound in [18] is a powerful tool under its specific conditions, extrapolating it beyond these conditions is a shaky practice.
@@
$\mathrm{MSE}(y_{0})\coloneqq\underset{\zeta=(\zeta_{0},\dots,\zeta_{N-1})}{\inf}\ \mathbb{E}\big{[}\big{|}Y_{N}^{h,y_{0},\zeta}-g(X_{N}^{h,y_{0},\zeta})\big{|}^{2}\Big{]}.$ (6)
This is the MSE with $y_{0}$ fixed and only $\zeta$ being optimized. If joint optimization of $y_{0}$ and $\zeta$, as in the deep BSDE method, is supposed to yield a good approximation of the FBSDE, then by necessity the minimum of $y_{0}\mapsto\mathrm{MSE}(y_{0})$ must be close to the true $Y_{0}$ of the FBSDE. For an LQ problem, detailed in Section 4.1.2 below, Figure 1 shows that the optimization landscape $y_{0}\mapsto\mathrm{MSE}(y_{0})$ instead has a minimum $y_{0}^{*}$ very far from the true $Y_{0}$. Figure 2 shows the exact $Y$ and the output of the deep FBSDE method, pathwise and in mean. We note that although the initial condition is significantly incorrect, the terminal condition is well approximated, i.e., $g(X_{N}^{h,y_{0},\zeta})-Y_{N}^{h,y_{0},\zeta}$ is small both on the pathwise level and in the mean-squared sense. It should be noted that the numerical algorithm has converged, as the loss function is no longer decreasing and $y_{0}$ remains constant (not shown here but Figure 6 qualitatively shows the same behavior for a different problem). Moreover, the algorithm is consistent, with repeated runs yielding similar values for the loss function and $y_{0}$. The reader might wonder if a smaller time step $h$ in the approximation would improve approximation of $Y_{0}$. Figure 5 indicates a negative answer to this valid thought by showing non-convergence of our implementation of the deep BSDE method (after thorough hyperparameter optimization) (the case $K=1$ in that figure).
Figure 1: MSE plotted for different values of $y_{0}$ for the LQ-problem of Section 4.1.2.
@@
Figure 2: The almost analytical $Y$-process and the approximation obtained by using the deep BSDE method. Left: One representative path. Right: A sample mean of size $2^{12}$.
@@ 3.2 The deep multi-FBSDE method
$\begin{dcases}X_{t}^{\psi}=x_{0}+\int_{0}^{t}\big{(}b(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})-\psi(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})\big{)}\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\psi})\text{d}W_{s},\\ Y_{t}^{\psi}=g(X_{T}^{\psi})+\int_{t}^{T}\big{(}f(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi})+\big{\langle}\psi(s,X_{s}^{\psi},Y_{s}^{\psi},Z_{s}^{\psi}),Z_{s}^{\psi}\big{\rangle}\big{)}\text{d}s\\ \quad\quad\ -\int_{t}^{T}\big{\langle}Z_{s}^{\psi},\sigma(s,X_{s}^{\psi})\text{d}W_{s}\big{\rangle},\quad t\in[0,T].\end{dcases}$ (7)
@@
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{T}^{y_{0},\zeta,\psi}-g(X_{T}^{y_{0},\zeta,\psi})\big{|}^{2},\quad\text{where}\\ X_{t}^{y_{0},\zeta,\psi}=x_{0}+\int_{0}^{t}b^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta,\psi}=y_{0}-\int_{0}^{t}f^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\big{\langle}Z^{y_{0},\zeta,\psi}_{s},\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s}\big{\rangle},\\ Z^{y_{0},\zeta,\psi}_{t}=\zeta(t,X_{t}^{y_{0},\zeta,\psi}),\quad t\in[0,T].\end{dcases}$ (8)
@@
$\begin{dcases}\underset{y_{0},\zeta}{\mathrm{minimize}}\ \sum_{k=1}^{K}\mathbb{E}\big{|}Y_{T}^{y_{0},\zeta,\psi_{k}}-g(X_{T}^{y_{0},\zeta,\psi_{k}})\big{|}^{2},\quad\text{where for }\psi\in\{\psi_{1},\dots,\psi_{K}\}\\ X_{t}^{y_{0},\zeta,\psi}=x_{0}+\int_{0}^{t}b^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s},\\ Y_{t}^{y_{0},\zeta,\psi}=y_{0}-\int_{0}^{t}f^{\psi}(s,X_{s}^{y_{0},\zeta,\psi},Y_{s}^{y_{0},\zeta,\psi},Z^{y_{0},\zeta,\psi}_{s})\text{d}s+\int_{0}^{t}\big{\langle}Z^{y_{0},\zeta,\psi}_{s},\sigma(s,X_{s}^{y_{0},\zeta,\psi})\text{d}W_{s}\big{\rangle},\\ Z^{y_{0},\zeta,\psi}_{t}=\zeta(t,X_{t}^{y_{0},\zeta,\psi}),\quad t\in[0,T].\end{dcases}$ (9)
@@
$\begin{dcases}\underset{y_{0},\zeta_{0},\ldots,\zeta_{N-1}}{\mathrm{minimize}}\ \sum_{k=1}^{K}\mathbb{E}\big{|}Y_{N}^{h,y_{0},\zeta,\psi_{k}}-g(X_{N}^{h,y_{0},\zeta,\psi_{k}})\big{|}^{2},\quad\text{where for }\psi\in\{\psi_{1},\dots,\psi_{K}\}\\ X_{n+1}^{h,y_{0},\zeta,\psi}=X_{n}^{h,y_{0},\zeta,\psi}+b^{\psi}\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi},Y_{n}^{h,y_{0},\zeta,\psi},Z_{n}^{h,y_{0},\zeta,\psi}\big{)}h+\sigma(t_{n},X_{n}^{h,y_{0},\zeta,\psi})\Delta W_{n},\\ Y_{n+1}^{h,y_{0},\zeta,\psi}=Y_{n}^{h,y_{0},\zeta,\psi}-f^{\psi}\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi},Y_{n}^{h,y_{0},\zeta,\psi},Z_{n}^{h,y_{0},\zeta,\psi}\big{)}h+\big{\langle}Z_{n}^{h,y_{0},\zeta,\psi},\sigma\big{(}t_{n},X_{n}^{h,y_{0},\zeta,\psi}\big{)}\Delta W_{n}\big{\rangle},\\ Z_{n}^{h,y_{0},\zeta,\psi}=\zeta_{n}(X_{n}^{h,y_{0},\zeta,\psi}\big{)},\ X_{0}^{h,y_{0},\zeta,\psi}=x_{0},\ Y_{0}^{h,y_{0},\zeta,\psi}=y_{0},\\ \text{for }n\in\{0,1,\ldots,N-1\}.\end{dcases}$ (10)
@@
$\begin{dcases}\underset{\zeta_{0},\ldots,\zeta_{N-1}}{\mathrm{minimize}}\ \mathbb{E}\big{|}Y_{N}^{h,\zeta}-g(X_{N}^{h,\zeta})\big{|}^{2},\quad\text{where}\\ X_{n+1}^{h,\zeta}=X_{n}^{h,\zeta}+b\big{(}t_{n},X_{n}^{h},Y_{n}^{h,\zeta},Z_{n}^{h}\big{)}h+\sigma(t_{n},X_{n}^{h,\zeta})\Delta W_{n},\\ Y_{n+1}^{h,\zeta}=Y_{n}^{h,\zeta}-f\big{(}t_{n},X_{n}^{h,\zeta},Y_{n}^{h,\zeta},Z_{n}^{h,\zeta}\big{)}h+\big{\langle}Z_{n}^{h,\zeta},\sigma\big{(}t_{n},X_{n}^{h,\zeta}\big{)}\Delta W_{n}\big{\rangle},\\ Z_{n}^{h,\zeta}=\zeta_{n}\big{(}X_{n}^{h,\zeta}\big{)},\ Y_{0}^{h,\zeta}=y_{0}^{*},\quad\text{for }n\in\{0,1,\ldots,N-1\}.\end{dcases}$ (11)